# Primordial Black Holes as Seeds for Extremely Overmassive AGN Observed by JWST

### v1

@@ 4 Limitation and Caveats
  • Similar to Zhang2025PBHStar, our physical treatment of the PBH seed remains idealized. A single PBH seed is placed in an isolated box, neglecting potentially important effects such as an extended PBH mass function, PBH clustering, or mergers with forming galaxies. A broader PBH mass spectrum could alter small-scale structure formation (e.g., Delos2024JCAP...12..005D), while clustering or early mergers may enhance gas inflow, BH growth, and star formation by supplying additional cold material (see a review of dual and binary SMBHs in DeRosaSMBHB2019NewAR..8601525D). Future work will explore a wider range of PBH masses, spatial configurations, and accretion prescriptions to quantify their impact on early SMBH–galaxy coevolution.
@@
  • A related limitation concerns our modeling of the dark matter component. In the present simulations, the PBH grows within collisionless PDM that does not undergo annihilation, self-interactions, or capture by the central PBH. As a result, the halo develops and maintains a steep inner cusp in PBH_SF_M5e7_fd005 (as shown in Figure 1). This profile is qualitatively consistent with analytic expectations for the DM profile around a central compact object under adiabatic dynamical evolution (e.g., Merritt2004PhRvL..92t1304M). In mixed dark matter scenarios, however, the long-term evolution of the cusp may differ: annihilation can flatten the inner density profile, while PBH–DM capture or dynamical heating can remove dark matter from the center. In our simulations, the formation and evolution of PDM density spikes are related to both the density perturbations developed around the PBH during the radiation-dominated epoch (see e.g., Boucenna2018PBHWIMP; Carr2021MNRAS.506.3648C, also captured by the initial conditions of our simulations) and the subsequent in-fall of gas towards the center throughout the cosmic history. The interplay of these processes is still not fully understood. Capturing these processes and comparing to analytical predictions requires a more sophisticated treatment of dark matter microphysics, its interaction with baryons, and BH accretion feedback, which we leave for future work.
@@
  • Modeling stellar feedback in PBH-seeded systems is intrinsically complex, involving the interplay of radiative heating, SN explosions, and black-hole–driven outflows. In the present implementation, SN feedback is treated in a subgrid manner: each stellar particle releases metals and injects thermal energy into its surroundings at the end of its life. Metals are distributed uniformly in SN bubbles. However, this treatment neglects spatial inhomogeneities and small-scale turbulent mixing during enrichment (e.g., Ritter_2015). As a result, it may underestimate local metallicity variations by overestimating the efficiency of chemical homogenization within the ISM (e.g., Sarmento2022ApJ...935..174S; Steinwandel2025ApJ...991...16S). Incorporating explicit metal-diffusion models and (sub-grid) turbulence-driven mixing prescriptions would yield a more physically accurate representation of early chemical enrichment in PBH-seeded galaxies.
@@
  • Our simulations assume BH feedback is purely thermal. However, at later stages, effects such as radiation pressure, mechanical winds, or jet-driven feedback may become important (see the review by Inayoshi:2020). These processes can transport metal-enriched gas into the circumgalactic medium more efficiently. A more complete treatment that couples radiative transfer, kinetic feedback, and a multiphase ISM will therefore be essential for capturing the full range of SMBH-driven regulation in early galaxies.
5 Conclusion
@@
The resulting system—whether observed immediately prior to star formation or shortly after the initial burst—naturally exhibits both low metallicity and an extreme black-hole-to-stellar mass ratio, reproducing the key properties of Abell 2744–QSO1. Moreover, the predicted PBH accretion rates, primarily regulated by BH feedback in our simulations, are consistent with the value inferred from observations. Early dominance of BH feedback thus provides a compelling physical explanation for the compact, metal-poor, BH-dominated systems uncovered by JWST.
Distinguishing observationally between a dense, star-free gas cocoon and a compact stellar cluster surrounding a primordial seed will require detailed modeling of the emergent spectral energy distribution (SED), including signatures of massive star formation, nebular line emission, and reprocessed radiation. In practice, however, these phases may be difficult to separate, since the central BH is enshrouded by dense metal-poor gas in both scenarios and the stellar component in our model is subdominant. It is also plausible that JWST may capture such systems during the brief, bursty star-formation episodes associated with enhanced BH accretion, as a natural outcome of PBH-regulated growth. Future work will incorporate radiative-transfer tools such as CLOUDY (CLOUDY2023RMxAA..59..327C) to generate self-consistent SEDs and mock observations, allowing more robust predictions for observational diagnostics. Additional simulations in cosmological environments that permit PBH clustering, interactions, and early mergers will further enable a more complete characterization of the diversity of early SMBH–host galaxy coevolution pathways.
We acknowledge fruitful discussions with Priyamvada Natarajan and Mike Boylan-Kolchin. The authors acknowledge the Texas Advanced Computing Center (TACC) for providing HPC resources under allocation AST23026. BL gratefully acknowledges the funding of the Royal Society University Research Fellowship and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).

### v2

@@ IV Limitation and Caveats
  • Similar to Zhang et al. (2025b), our physical treatment of the PBH seed remains idealized. A single PBH seed is placed in an isolated box, neglecting potentially important effects such as an extended PBH mass function, PBH clustering, or mergers with forming galaxies. A broader PBH mass spectrum could alter small-scale structure formation (e.g., Delos et al., 2024), while clustering or early mergers may enhance gas inflow, BH growth, and star formation by supplying additional cold material (see a review of dual and binary SMBHs in De Rosa et al., 2019). Future work will explore a wider range of PBH masses, spatial configurations, simulation box sizes, and accretion prescriptions to quantify their impact on early SMBH–galaxy coevolution.
@@
  • A related limitation concerns our modeling of the dark matter component. In the present simulations, the PBH grows within collisionless PDM that does not undergo annihilation, self-interactions, or capture by the central PBH. As a result, the halo develops and maintains a steep inner cusp in PBH_SF_M5e7_fd005 (as shown in Figure 1). This profile is qualitatively consistent with analytic expectations for the DM profile around a central compact object under adiabatic dynamical evolution (e.g., Merritt, 2004). In mixed dark matter scenarios, however, the long-term evolution of the cusp may differ: annihilation can flatten the inner density profile, while PBH–DM capture or dynamical heating can remove dark matter from the center. In our simulations, the formation and evolution of PDM density spikes are related to both the density perturbations developed around the PBH during the radiation-dominated epoch (see, e.g., Boucenna et al., 2018; Carr et al., 2021c, also captured by the initial conditions of our simulations) and the subsequent in-fall of gas towards the center throughout the cosmic history. The interplay of these processes is still not fully understood. Capturing these processes and comparing to analytical predictions requires a more sophisticated treatment of dark matter microphysics, its interaction with baryons, and BH accretion feedback, which we leave for future work.
@@
  • Modeling stellar feedback in PBH-seeded systems is intrinsically complex, involving the interplay of radiative heating, SN explosions, and black-hole–driven outflows. In the present implementation, SN feedback is treated in a subgrid manner: each stellar particle releases metals and injects thermal energy into its surroundings at the end of its life. Metals are distributed uniformly in SN bubbles. However, this treatment neglects spatial inhomogeneities and small-scale turbulent mixing during enrichment (e.g., Ritter et al., 2015). As a result, it may underestimate local metallicity variations by overestimating the efficiency of chemical homogenization within the ISM (e.g., Sarmento and Scannapieco, 2022; Steinwandel et al., 2025). Incorporating explicit metal-diffusion models and (sub-grid) turbulence-driven mixing prescriptions would yield a more physically accurate representation of early chemical enrichment in PBH-seeded galaxies.
@@
  • Our simulations assume BH feedback is purely thermal. However, at later stages, effects such as radiation pressure, mechanical winds, or jet-driven feedback may become important (see the review by Inayoshi et al., 2020). These processes can transport metal-enriched gas into the circumgalactic medium more efficiently. A more complete treatment that couples radiative transfer, kinetic feedback, and a multiphase ISM will therefore be essential for capturing the full range of SMBH-driven regulation in early galaxies.
In this work, we focus on a proof-of-concept demonstration showing that a massive PBH seed can reproduce the observed properties of Abell 2744–QSO1. If one assumes that the entire JWST “little red dot” population originated from a similar PBH-seeded pathway, then using a measured number density of $\sim 4\times 10^{-4}\,\mathrm{cMpc}^{-3}$ at $6.5<z<8.5$ for a rough estimate (e.g., see table 2 in Kokorev et al., 2024) would imply an upper limit on the density fraction of $5\times 10^{7}\ \mathrm{M}_{\odot}$ PBH seeds of $f_{\rm PBH}\lesssim 3.2\times 10^{-7}$, much smaller than the existing constraints from dynamical effects and also below typical CMB heating/early-accretion limits (see, e.g., Sec. III.E of Carr et al., 2021b)  [^12] For $30\,M_{\odot}<M_{\rm PBH}\lesssim 10^{4}\,M_{\odot}$, a commonly used approximate scaling for the CMB heating/early-accretion limit is $f_{\rm PBH}\lesssim(M_{\rm PBH}/30\,M_{\odot})^{-2}$, with a high-mass flattening at the level of $f_{\rm PBH}\lesssim 10^{-5}$ under Eddington-limited accretion (see related treatments in Ricotti et al., 2008; Ali-Haïmoud and Kamionkowski, 2017). On the other hand, CMB $\mu$-distortion constraints tightly limit PBHs at this mass scale (Nakama et al., 2018; Hooper et al., 2024; De Luca et al., 2025), but these bounds assume Gaussian primordial fluctuations; non-Gaussianity or alternative formation scenarios can substantially weaken them (see, e.g., section II in Escrivà et al., 2024).. We note, however, that several recent studies have proposed alternative formation channels for LRDs and QSO1-like systems(e.g., Lu et al. (2024); Qin et al. (2025); Jiang et al. (2026); De Luca et al. (2025), and see section 7 in (Maiolino et al., 2025)). Like our study, most of these works remain at the proof-of-concept stage and lack the follow-up numerical modeling required for a quantitative comparison across scenarios. A systematic assessment of competing pathways—spanning population statistics, chemical enrichment histories, environmental/clustering signatures, and multi-wavelength diagnostics—will therefore require substantial future work.
V Conclusion
@@
The resulting system—whether observed immediately prior to star formation or shortly after the initial burst—naturally exhibits both low metallicity and an extreme black-hole-to-stellar mass ratio, reproducing the key properties of Abell 2744–QSO1 (as shown in Table 2). Moreover, the predicted PBH accretion rates, primarily regulated by BH feedback in our simulations, are consistent with the value inferred from observations. Early dominance of BH feedback thus provides a compelling physical explanation for the compact, metal-poor, BH-dominated systems uncovered by JWST.
Table 2: Comparison of key observed properties of Abell 2744–QSO1 (adapted from Table 2 in Maiolino et al., 2025) and the corresponding values from our PBH-seeded simulation (the PBH_SF_M5e7_fd005 run) snapshot at $z\simeq 7$.
Quantity Observation (QSO1) Simulation ($z\simeq 7$)
BH Mass, $M_{\rm BH}$($\mathrm{M}_{\odot}$]) $\sim(2.5$$10)\times 10^{7}$ $\sim 6\times 10^{7}$
Stellar Mass, $M_{\star}$($\mathrm{M}_{\odot}$]) $\lesssim 2\times 10^{7}$ (dynamics), $\sim 10^{6}$ (optical and UV) $\sim 7.7\times 10^{5}$
$M_{\rm BH}/M_{\star}$ $\gtrsim 2$ (model dependent) $\sim 80$
Metallicity ($R<150$ pc; $Z/Z_{\odot}$]) $\sim 10^{-2.23}-10^{-2.44}$ $\sim 10^{-2.28}$
Metallicity ($150$ pc <R<300<R<300 pc; $Z/Z_{\odot}$]) $\lesssim 10^{-2.41}$ $\sim 10^{-2.31}$
Eddington Ratio, $f_{\rm Edd}$ $\sim 0.01$–0.03 $\sim 0.02$
Distinguishing observationally between a dense, star-free gas cocoon and a compact stellar cluster surrounding a primordial seed will require detailed modeling of the emergent spectral energy distribution (SED), including signatures of massive star formation, nebular line emission, and reprocessed radiation. In practice, however, these phases may be difficult to separate, since the central BH is enshrouded by dense metal-poor gas in both scenarios and the stellar component in our model is subdominant. It is also plausible that JWST may capture such systems during the brief, bursty star-formation episodes associated with enhanced BH accretion, as a natural outcome of PBH-regulated growth. Future work will incorporate radiative-transfer tools such as CLOUDY (Chatzikos et al., 2023) to generate self-consistent SEDs and mock observations, allowing more robust predictions for observational diagnostics. Additional simulations in cosmological environments that permit PBH clustering, interactions, and early mergers will further enable a more complete characterization of the diversity of early SMBH–host galaxy coevolution pathways. In parallel, future JWST ultra–deep surveys will be crucial for establishing a statistical census of LRDs. By increasing the number of spatially resolved, high–signal–to–noise detections of low–metallicity, BH–dominated systems, such surveys will enable population–level diagnostics such as number–density estimates, the distribution of $M_{\rm BH}/M_{\star}$, spatial metallicity gradients, and the fraction of sources exhibiting sub–Eddington accretion. These trends will provide key leverage for distinguishing PBH seeds from alternative formation pathways (see e.g., Qin et al., 2025; Jeon et al., 2025b), which predict different duty cycles, enrichment histories, and host–environment signatures.
We acknowledge fruitful discussions with Priyamvada Natarajan, Fangzhou Jiang, Kohei Inayoshi, and Barmak Shams Es Haghi and within Mike Boylan-Kolchin’s group. The authors acknowledge the Texas Advanced Computing Center (TACC) for providing HPC resources under allocation AST23026. BL gratefully acknowledges the funding of the Royal Society University Research Fellowship and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).